Pls. help me solve this problem.
Ana invested 50,000. Part of it she invested in a mine stock from which she hoped to receive a 20 percent return per yr. The rest she invested in a bank stock which was paying 6 percent per year. If she received $400 more the first year from the bank stock than from the mining stock, how much did she invest in each stock.

Answer by mananth(16946)

Ana invested 50,000. Part of it she invested in a mine stock from which she hoped to receive a 20 percent return per yr. The rest she invested in a bank stock which was paying 6 percent per year. If she received $400 more the first year from the bank stock than from the mining stock, how much did she invest in each stock.
...
Ana invested 50,000.
mine stock $ x------- 20 %
bank stock $ (50000-x)-----6 %
If she received $400 more the first year from the bank stock than from the mining stock, how much did she invest in each stock.
Bank stock interest - mine stock interest = 400
(50,000-x)*0.06 - 0.20x = 400
multiply by 100
(50,000-x)*6-20x=400*100
300000-6x-20x=40000
-26x=-260,000
/-26
x= 10000
$10,000 money invested in mine stock
$40,000 in bank stock